### v1

# Topology of Fermi seas and geometry of their boundaries for free particles in one and two-dimensional lattices

Topology of Fermi seas and geometry of their boundaries for free particles in one and two-dimensional lattices
Guillermo R. Zemba
Facultad de Ingeniería y Ciencias Agrarias, Pontificia Universidad Católica Argentina,
Av. Alicia Moreau de Justo 1500,(C1107AAZ) Buenos Aires, Argentina
and
Departamento de Física Teórica de Interacciones Fundamentales y Sistemas Complejos, Laboratorio Tandar,
Comisión Nacional de Energía Atómica,
Av.Libertador 8250,(C1429BNP) Buenos Aires, Argentina
Free gases of spinless fermions moving on a geometric background with lattice symmetries are considered. Their Fermi seas and corresponding boundaries may be classified according to their topological properties at zero temperature. This is accomplished by considering the flat orbifolds ${\mathbb{R}}^{d}/\Gamma$, with $\Gamma$ being the crystallographic group of symmetry in $d$-dimensional momentum space. For $d=1$, there are 2 topological classes: a circumference, corresponding to an insulator and an interval, identified as a conductor. For $d=2$, the number of topological classes extends to 17: there are 8 with the topology of a disk identified as conductors and 4 corresponding to a 2-sphere matching insulators, both sets eventually including finite numbers of conical singularities and reflection corners at the boundaries. The rest of the listing includes single cases corresponding to insulators (2-torus, real projective plane, Klein bottle) and conductors (annulus, Möbius strip). Physical interpretations of the singularities are provided, as well as examples that fit within this listing. Given the topological nature of this classification, its results are expected to be robust against small perturbative interactions.
1 Introduction
In the last decades, great progress has been made in the theoretical understanding of the properties of materials from their microscopic structure. One of the main devices for this progress is provided by band theory [1, 2, 3], which offers a potent framework for understanding and predicting the electronic properties of solids by describing the possible energy levels of electrons moving in a lattice structure [4, 5]. According to the band structures arising from these studies, materials may be classified as conductors, insulators, or semiconductors. Moreover, the modification of material properties by a change in its microscopic structure is within the reach of these methods. On a given material, one of the most relevant consequences of band structure studies is the determination of its Fermi sea and Fermi surface. Since lattices in 1 and 2 spatial dimensions will be considered in this study, the term Fermi boundary instead of Fermi surface will be employed, for the sake of clarity. For a given energy function, the Fermi boundary separates occupied (low energy) from unoccupied electron states in momentum space, for a material at zero temperature. This definition generalizes the paradigm of a free fermion gas to the case of a gas moving on a spatial lattice background with no further structure beyond its geometry. Moreover, the shape and topology of the Fermi sea and its boundary encode crucial information about the material properties in a compact manner [6], for instance, their electrical behavior. The predicted momentum space properties of the Fermi sea are also confirmed by experimental techniques like angle-resolved photoemission spectroscopy (ARPES) and the observation of the De Haas-van Alphen effect [7]. One the other hand, band studies usually demand extensive use of computational resources. That is, extensive detailed information provided as an input yields extensive detailed information as an output.
On the other hand, effective field theories (EFT) provide alternative, mostly analytic methods to study the motion of fermions in crystal backgrounds that are valid in the long-distance, low-energy realm of a system in the thermodynamic limit (large number of non-relativistic particles at constant spatial density) [8, 9]. These methods require much less detailed microscopic information and therefore provide accordingly a commensurate output of information. The key features here is that the outcome consists of universal information, independent of the microscopic details and requires only simple computation tools. For free particles, the input information is basically provided by the symmetries of the microscopic lattice. The Fermi sea and energy are defined by the free hamiltonian describing the dynamics of the particles, i.e., possessing kinetic energy only. This starting point based on the free theory is adopted following the EFT method following the Wilson Renormalization Group ideas [8]. It should be, therefore, feasible to determine the universal properties of the Fermi seas and their boundaries just from the symmetry contents of the lattice. These should be the topological and geometrical structures of the Fermi seas and boundaries, leaving aside their detailed quantitative properties. The inclusion of interactions among the particles and with a potential background (e.g., provided by a crystal of ions) is expected to leave invariant the topological structures of the Fermi seas and their boundaries.
The free fermion gas picture describes the propagation of $N$ non-relativistic, non-interacting fermion particles in $d$-dimensional Euclidean space, often considered confined to a large volume, in the limit in which the spatial density of the particles $n(d)=N/V(d)$ is constant, i.e., $N\to\infty$, $V(d)\to\infty$, where $V(d)$ the $d$-dimensional volume. This picture may be extended to the case of the same type of particles propagating in the presence of a background lattice [9], considered just as a modification of the free geometry and not as an interacting background potential. The momentum space geometry, defined as the available space to be occupied by free fermionic states, is modified from ${\mathbb{R}}^{d}$ to the quotient space ${\mathbb{R}}^{d}/\Gamma$, where $\Gamma$ is the crystallography group of symmetry of the lattice in $d$-dimensional momentum space. This topological space is known as an orbifold (see e.g., [10, 14]). One caveat regarding this statement is that only the compact connected orbifolds that are flat (e.g., in $d=2$ this means that their Euler characteristic is zero) should be considered, as they should locally resemble ${\mathbb{R}}^{d}$. This is because orbifolds could have point singularities that modify its spatial curvature.
The following two Sections present and discuss the $1d$ and $2d$ cases. The $3d$ case is outlined before the presentation of Concluding Remarks.
2 The $1d$ case
In $d=1$ a gas of non-relativistic spinless free fermions of mass $m$ is described by $N$ non-interacting particles moving on the space $\mathbb{R}$} with coordinate $x$ and macroscopic size $L$. Their hamiltonian is given by the kinetic energy $H=p^{2}/(2m)$ and the solutions to the Schrödinger equation are free plane wave functions $e^{ipx}$ ($\hbar=1$) with quantized momenta , with $n$ being an integer or half-integer, according to the chosen boundary conditions at $x=0$ and $x=L$ . The dispersion relation is $E(p)=p^{2}/(2m)$, with $E(p)$ being the energy. In the thermodynamic limit $N\to\infty$, $L\to\infty$ with $n(1)=N/L$ constant, the Fermi sea at zero temperature is the configuration obtained by filling all the lowest energy $N$ states in momentum space so as to form a segment with boundary points -$p_{F}$ and $p_{F}$, where $p_{F}=\pi n(1)$ is the Fermi momentum. Therefore, the Fermi sea is an interval in momentum space and its Fermi boundary is composed by two points [1]. Note that the given geometry of the space of possible momenta is $\mathbb{R}$.
The above picture may be modified by the introduction of a background spatial lattice, which only changes the geometrical properties of free propagation (i.e., no interaction with a crystal implied). In $d=1$, there is only type of Bravais lattice consisting of points equally spaced along the $\mathbb{R}$ line, with spacing $a$, which is a microscopic scale. The reciprocal or momentum lattice is obtained by Fourier transformation of the spatial (direct) one and found to be isomorphic to it, with spacing $2\pi/a$, which is a large scale. The free hamiltonian is replaced by a periodic version of it, the hoping hamiltonian:
$H=-t[\cos(pa)-1]\ ,$ (1)
where the parameter $t$ is related to the mass of the particles by $t=1/(ma^{2})$ after considering the limit of small momentum of (1). Note that (1) incorporates the reflection ($P$) $\mathbb{Z}_{2}$ symmetry: $p\to-p$. The solutions of the Schrödinger equation are now given by Bloch wave functions $u(x)e^{ikx}$, with $u(x+2\pi/a)=u(x)$ [1, 2], and the dispersion relation becomes $E(p)=-t[\cos(pa)-1]$. It is convenient to define the number of spatial cells $N_{c}=L/a$ and the number of particles per unit cell $\nu=Z=N/N_{c}$ so that $n(1)=\nu/a$. The geometry of the momentum space states is, therefore, realized as a ring of radius $1/a$. Given the $2\pi/a$ periodicity of momenta, these belong to equivalent regions identified as Brioullin zones (BZ). Now the geometry of the space of available momenta is the circumference $S^{1}$ , representing the ring. One could imagine filling the lowest momentum states with free fermions starting from the circumference angle parametrization $\theta=0$ up to a value $\theta$ for $N/2$ particles with positive momentum and -$\theta$ for the remaining $N/2$ particles with negative momentum, due to the $T$ symmetry. Therefore, the Fermi sea geometry is the interval $p_{F},p_{F}]$ and the Fermi boundary consists of two isolated points at $\pm p_{F}$ with $2p_{F}=(2\pi/L)N=(2\pi/a)\nu$. The topology of the Fermi sea is, instead, given by the fundamental domain of momenta, that is the subset of points of the Fermi sea that are identified by the $P$ symmetry. It is given by the arc $p>0$, i.e., the irreducible wedge of the $1d$ lattice, noting that the above picture may be folded about the abscissa axis. The space of free independent momentum states is, therefore, the interval $(0,p_{F}]$ (the zero momentum state should be excluded due to the uncertainty principle given the finite size of the system). This folding is known as the orbifolding of the $S^{1}$ manifold by the $P$ ($\mathbb{Z}_{2}$) symmetry, i.e., the quotient space $S^{1}/\mathbb{Z}_{2}$. In conclusion, the generic topology of the Fermi sea is that of the interval. The physical systems may be identified as conductors, in the sense that there are available particle-hole low-lying excitations about each of the Fermi points. There is, however, a special configuration reached when the filled states in the momentum ring cover it up exactly, that is, when twice the Fermi momentum equals the lattice spacing $2\pi/a$, meaning that both Fermi points hit the boundaries of the first BZ: $p_{F}=\pi n(1)=\pi/a$ , i.e., $\nu=1$. Due to the ring topology, both Fermi points coalesce (are identified) so that the Fermi sea becomes a circumference with no Fermi boundary. Therefore, no low-lying excitations are available and the systems are identified as insulators (more properly, they could be termed as geometric insulators. The topology of the Fermi sea is now given by the circumference $S^{1}$.
The same results could have been obtained from a purely mathematical analysis. In $d=1$ there are two possible space groups $\Gamma$: $\mathbb{Z}$ and $\mathbb{D}_{\infty}$ (see Table 1) [10]. These arise from identification of the isometries of $\mathbb{R}$ that act as generators of the group. The orbifolds (quotient spaces) $\mathbb{R}/\Gamma$ are two: the circumference and the interval.
$\mathbb{Z}$ $\mathbb{R}/\mathbb{Z}$ $S^{1}$ circumference
$\mathbb{D}_{\infty}$ $\mathbb{R}/\mathbb{D}_{\infty}$ $S^{1}/\mathbb{Z}_{2}$ interval
Table 1: Space groups $\Gamma$ for $d=1$ are listed in the first column. The second stipulates the corresponding orbifold. The third and fourth indicate their topologies.
Is is worthwhile to mention that the space group $\mathbb{D}_{\infty}$ can be defined equally well in terms of two $\pi$-rotations $\mathbb{D}_{\infty}\simeq\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. The classification of $d=1$ momentum space orbifolds, therefore, yields the same resulting topologies of Fermi seas as the previous analysis based on the simple physical picture.
This mathematical analysis may be generalized to higher dimensional lattices (in particular, to $d=2$ and $d=3$) by promoting as a guiding principle the construction of all momentum orbifold spaces obtained from $\mathbb{R}^{d}$ and $\Gamma$, where $\Gamma$ denotes the crystallographic symmetry group of the lattice, which define the available spaces to be filled by free fermions.
3 The $2d$ case
In $d=2$ a gas of non-relativistic spinless free fermions of mass $m$ is described by $N$ non-interacting particles moving on the space ${\mathbb{R}}^{2}$ with coordinates $(x,y)$. For simplicity, a macroscopic size of $L^{2}$ is considered. Their free hamiltonian is given by $H=\left(p_{x}^{2}+p_{y}^{2}\right)/(2m)$. The construction parallels that of the $d=1$ case, with a $2d$ square lattice of momenta $(p_{x},p_{y})=(2\pi n_{x}/L,2\pi n_{y}/L)$ with $n_{x}$, $n_{y}$ integers. Now the geometry of the Fermi sea at zero temperature is a circle of radius $p_{F}$, due to the rotational invariance of the hamiltonian [1]. It is given by $p_{F}=2\sqrt{\pi n(2)}$, where $n(2)=N/L^{2}=\nu/a^{2}$ taking $L^{2}=N_{c}a^{2}$ and $\nu=Z=N/N_{c}$. A more detailed analysis of the topology of the Fermi sea concludes that it is given by the annulus $A$, which is a disk $D^{2}$ with a puncture (small hole) at the origin, a consequence of excluding the zero momentum state at the origin in the polar coordinates implied by the rotational symmetry about the origin. This is a flat manifold, with zero Euler characteristic (a measure of the curvature). Note that the given geometry of the space of possible momenta is the equally flat $\mathbb{R}^{2}$.
A simple extension of the $1d$ example is defined by free fermions on a spatial square lattice of side $a$. The momentum space lattice is given by a square of side $2\pi/a$. The free hoping hamiltonian that is consistent with this geometry is given by:
$H=-t[\cos(p_{x}a)+\cos(p_{y}a)-2]\ .$ (2)
This hamiltonian does not have rotational invariance in momentum space, but rather the discrete symmetries $p_{x}\leftrightarrow p_{y}$, $p_{x}\leftrightarrow-p_{x}$ and $p_{y}\leftrightarrow-p_{y}$. Now the Fermi sea has the shape of a diamond in momentum space. The first BZ is given by the square of side intervals $(-\pi/a,\pi/a)$ in both the $p_{x}$ and $p_{y}$ axes. The Fermi momentum $p_{F}$ is given by half the length of the side of the diamond. These diamond shapes have four corners at an angle of $\pi/2$ that may be regarded as Van Hove singularities, such that the Fermi velocity (the gradient of the energy with respect to the momentum) vanishes there [8]. This is the generic picture for $\nu<1$, which describe conductors generically. For example, for $\nu=1/2$, the vertices of the diamond are located at $(\pm\pi/a,\pm\pi/a)$. Thus, the geometry of the Fermi sea for $\nu<1$ is that of a diamond, and generically describes a conductor. Its topology is that of a disk, with singular boundary. As in the $d=1$ case, a special configuration exists for $\nu=1$. The Fermi sea extends beyond the first BZ, with vertices of the diamond located at $\sqrt{2}(-\pi/s,\pi/a)$ and $p_{F}=\pi/a$. This means that the opposite sides of the Fermi sea can be mapped one into another by the addition of a vector proportional to $(\pm 2\pi/a,\pm 2\pi/a)$, meaning that they are equivalent in the lattice and, therefore, identified (the so called nesting mechanism). The topology of the Fermi sea becomes that of a $2d$ torus $T^{2}$, with no boundary and interpreted as an insulator. This is just one example out of the many that may be explicitly considered in $2d$.
On the other hand, the mathematical analysis is more powerful and goes as follows. There are 5 types of $2d$ Bravais lattices: oblique (no restrictions on the angles or lengths of the unit cell vectors), rectangular ($\pi/2$ angles, unequal side lengths), centered rectangular ($\pi/2$ angles, unequal side lengths, centered unit cell), square ($\pi/2$ angles, equal side lengths) and hexagonal ($2\pi/3$ and $\pi/3$ angles, equal side lengths) [1, 2]. The reciprocal of each $2d$ Bravais lattice is also a $2d$ Bravais lattice, with the same symmetry as that of the direct one. The point groups describe the (rotational, reflection, etc.) symmetries of the lattice. In $2d$, there are 17 distinct crystallographic space groups, also known as wallpaper groups. These groups are formed by combining Bravais lattices with different point group symmetries. The set of isometries comprises translations, reflections, $2\pi/n$-rotations with order $n=2,3,4,6$ and so-called glide-reflections, which are translations with a simultaneous mirror reflection. The crystallographic notation [12] is common in the physics literature for the space groups, whereas in mathematics, Conway’s orbifold notation [13] is more widely used. The classification of the 17 wallpaper groups is done by considering each Bravais lattice type and its possible point group symmetries at a time [16]. The list comprises:
  1. 1.
    15 primitive lattices:
    1. (a)
      2 parallelogram lattices: $\mathsf{p1},\mathsf{p2}$
    2. (b)
      5 rectangular lattices: $\mathsf{pm},\mathsf{pg},\mathsf{pmm},\mathsf{pmg},\mathsf{pgg}$
    3. (c)
      3 square lattices: $\mathsf{p4},\mathsf{p4g},\mathsf{p4m}$
    4. (d)
      3 trigonal lattices: $\mathsf{p3},\mathsf{p31m},\mathsf{p3m1}$
    5. (e)
      2 hexagonal lattices: $\mathsf{p6},\mathsf{p6m}$
  2. 2.
    2 centered lattices: $\mathsf{cm},\mathsf{cmm}$
[…]
As a consequence, there are 17 (compact and connected) orbifolds in $2d$ obtained as the quotient spaces $\mathbb{R}^{2}/\Gamma$, where $\Gamma$ is one of the space groups listed in Table 2 [10, 15].
$\mathsf{p2}$ $2222$ $\mathbb{R}^{2}/\mathsf{p2}$ 4-pillow $S^{2}$ 2,2,2,2
$\mathsf{p3}$ $333$ $\mathbb{R}^{2}/\mathsf{p3}$ 3-pillow 3,3,3
$\mathsf{p4}$ $442$ $\mathbb{R}^{2}/\mathsf{p4}$ 3-pillow 2,4,4
$\mathsf{p6}$ $632$ $\mathbb{R}^{2}/\mathsf{p6}$ 3-pillow 2,3,6
$\mathsf{pmm}$ 2222*$22$ $\mathbb{R}^{2}/\mathbb{D}_{\infty}^{2}$ rectangle $D^{2}$ 2,2,2,2
$\mathsf{cmm}$ 2222*22 $\mathbb{R}^{2}/\mathsf{cmm}$ triangle 2 2,2
$\mathsf{pmg}$ 2222* $\mathbb{R}^{2}/\mathsf{pmg}$ open 4-pillow 2,2 2
$\mathsf{p3m1}$ 333*333 $\mathbb{R}^{2}/\mathsf{p3m1}$ triangle 3,3,3
$\mathsf{p31m}$ 333*3 $\mathbb{R}^{2}/\mathsf{p31m}$ open 3-pillow 3 3
$\mathsf{p4m}$ 442*442 $\mathbb{R}^{2}/\mathsf{p4m}$ triangle 2,4,4
$\mathsf{p4g}$ 424*2 $\mathbb{R}^{2}/\mathsf{p4g}$ open 3-pillow 4 2
$\mathsf{p6m}$ 632*632 $\mathbb{R}^{2}/\mathsf{p6m}$ triangle 2,3,6
$\mathsf{p1}$ $\circ$ $\mathbb{R}^{2}/\mathbb{Z}^{2}$ torus $T^{2}$
$\mathsf{pg}$ $\times\times$ $\mathbb{R}^{2}/\mathsf{pg}$ Klein bottle $K$
$\mathsf{pgg}$ $22\times$ $\mathbb{R}^{2}/\mathsf{pgg}$ real projective plane $\mathbb{R}P^{2}$ 2,2
$\mathsf{pm}$ ** $\mathbb{R}^{2}/(\mathbb{Z}\times\mathbb{D}_{\infty})$ annulus AA
$\mathsf{cm}$ $\times$ $\mathbb{R}^{2}/\mathsf{cm}$ Möbius strip $M$
Table 2: The first two columns list the $2d$ space groups $\Gamma$ in crystallographic [12] as well as orbifold [13] notation. The following column displays the corresponding quotient spaces $\mathbb{R}^{2}/\Gamma$. The fourth and fifth column indicate the overall topology of the orbifold spaces. The sixth and seven list the order of conical points and the order of corner reflections, respectively.
The last three columns of Table 2 contain the topological classes of the Fermi seas and their boundaries. A schematic representation of every orbifold listed in Table 2 is provided in [17] in pages 79-80. There are 7 basic topologies: the 2-sphere ($S^{2}$), the 2-disk ($D^{2}$), the real projective plane $\mathbb{R}P^{2}$, the 2-torus ($T^{2}$), the Klein bottle ($K$), the annulus ($A$) and the Möbius strip ($M$). All of these are Euclidean planar spaces, such that their Euler characteristics is zero (flat orbifolds are known as parabolic). This property is automatic in the last 4 classes but not in the first 3. Indeed, the real projective plane needs a couple of conical singularities of order 2 to become flat. Moreover, the spaces $S^{2}$ and $D^{2}$ also require a prescribed number of isolated singularities so that the resulting space is flat: this results in 4 $S^{2}$ and 8 $D^{2}$ subclasses. These possibilities are determined from the general formula of the Euler characteristic for an orbifold $O$, which is given by (see, e.g., [15]):
$\chi(O)=\chi(M)-\sum_{i=1}^{p}\left(1-\frac{1}{m_{i}}\right)-\frac{1}{2}\sum_{j=1}^{q}\left(1-\frac{1}{n_{j}}\right)$ (3)
where $\chi(M)$ is the Euler characteristic of the smooth base manifold, $p$ are the cone points of order $m_{1},m_{2},\dots,m_{p}$ and $q$ the corner reflectors of order $n_{1},n_{2},\dots,n_{q}$.
A total of 7 orbifolds, on the listing in Table 2, are identified generically as insulators (no Fermi boundary): the 4 listed as $S^{2}$, the torus $T^{2}$, the real projective plane and the Klein bottle (the last 2 are non orientable in $3d$). The remaining 10 are identified generically as conductors (with a Fermi boundary): the 8 listed as $D^{2}$, the annulus and the Möbius strip (non orientable in $3d$). The reflection singularities located at the Fermi boundary can be identified as Van Hove singularities [2, 8]. The conical singularities located in the bulk of the Fermi sea could be associated to Dirac points [18], which means that there may be some selective, fine tuned conduction points with fixed Fermi velocities. Moreover, the appearance of Möbius-like Fermi surfaces has been studied in materials like twisted graphene with transition metal dichalcogenide (TMD) layers [19]. As a consistency test, it is noted that diamond shaped and the 2-torus Fermi seas found in the example of the square lattice are included in this classification. The first case corresponds to group $\mathsf{pmm}$ with symmetry $\mathbb{R}^{2}/\mathbb{D}_{\infty}^{2}$ and the second to the group $\mathsf{p1}$ with symmetry $\mathbb{R}^{2}/\mathbb{Z}^{2}$, which are the natural extensions of the $1d$ case, with symmetries $\mathbb{R}/\mathbb{D}_{\infty}$ (interval) and $\mathbb{R}/\mathbb{Z}$ (circumference), respectively.
The procedure of employing the mathematical classification of $2d$ flat orbifolds establishes not only the list of possible Fermi seas topologies and geometries of their boundaries but also the corresponding $2d$ lattice and its symmetry from which they are originated. However, the former information could be independently obtained by a more direct argument. Note that all the different subcases of the first two topological cases in Table 2, namely the 2-sphere with genus 2 and the disk with genus 1, may have been found by enumeration of all the possible singularities on these basic topologies, after imposition of the flatness condition. For the case of $S^{2}$, there are no boundaries and therefore no possible Van Hove singularities. The only possible ones are conical, of orders 2,3,4,6. Since the Euler characteristic of $S^{2}$ is 2 , the singularities should bring this number to 0 and therefore contribute to (3) with an amount -2. The only possible combinations are (3,3,3),(2,4,4),(2,3,6) and (2,2,2,2) matching the contents of Table 2, after use of (3). For the case of $D^{2}$, with the Euler characteristic 1, there are more possibilities given that the boundary is now the closed curve $S^{1}$. There are 4 possibilities that do not involve bulk singularities, and those are given by 3 triangles of internal angles $(\pi/3,\pi/3,\pi/3)$, $(\pi/2,\pi/4,\pi/4)$ and $(\pi/2,\pi/3,\pi/6)$ and 1 diamond of internal angles $(\pi/2,\pi/2,\pi/2,\pi/2)$. All of these configurations add up to a contribution of -1 to the Euler characteristic. Furthermore, there are 3 possibilities that combine boundary and bulk singularities: a droplet (tear) with a boundary angle of $\pi/2$ and a bulk singularity of order 4, another one analogous to the previous one with corresponding boundary angle $\pi/3$ and a bulk singularity of order 3 and a droplet with two boundary angles of $\pi/2$ and a bulk singularity of order 2. Finally, a configuration with only 2 bulk singularities of order 2.
A related classification of Fermi surface topological transitions in $2d$ based on Catastrophe theory is available in [20]. These transitions may be accounted for in the contents of Table 2 as topological operations linking pairs of the list. For instance, the $D^{2}$ configuration with two bulk singularities of order 2 could be brought to that of an annulus by merging the two conical singularities of order 2 into one of order 1 (that is, a puncture or another boundary with $S^{1}$ topology).
4 The $3d$ case: outline
In $d=3$, there are 14 unique Bravais lattices. These lattices are formed by combining 7 crystal systems with different centering types. The number of space groups is 230. A similar analysis to those of the $d=1$ and $d=2$ cases is feasible and postponed for further research.
5 Concluding remarks
Promoting the classification of the topology of Fermi seas by considering the orbifolds ${\mathbb{R}}^{d}/\Gamma$, where $\Gamma$ is the crystallography group of symmetry in $d$-dimensional momentum space, supported by a flatness condition, yields a listing that satisfies several consistency conditions and may be considered as an ordering principle for further research on this topic.
In $d=2$, the listing of possible topologies of the Fermi seas and the geometry of their boundaries may be obtained by a straightforward enumeration of all the possible conical and reflection singularities on surfaces with positive curvature (of genus $g=2,1$, that is, the 2-sphere and the disk) by imposing the flatness of the resulting orbifold.